# External charged debris in a flowing plasma : charge fluctuation induced complexity

### v2

@@ 2.4 Fluid model – FCT simulation
The FCT formalism requires the hydrodynamic equations Eqs.(3-6) to be put in the form of a generalized continuity equation,
$\frac{\partial f}{\partial t}=-\frac{\partial}{\partial x}(fv)+\frac{\partial s}{\partial x},$ (17)
where $f=(n_{i},v_{i})$ is the physical quantity to be solved, $(fv)$ is the corresponding flux, and $s$ is the source term. The external charged debris is entered through the Poisson equation Eq.(6) as before which is to be solved each time step along with Eq.(17). The external charged debris is modeled with the same profile as used in the PIC simulation, given by Eq.(12).
@@
$\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{m},\quad m_{{\rm FCT}}=6,$ (18)
@@
In the above expression, $R$ is the upper-triangular matrix as obtained from the QR decomposition, $p$ is the index of time step at which the decomposition is carried out, $N$ is total number of such steps, and $T$ is the physical time $T=\sum_{k=1}^{N}\Delta t_{k}$ with $\Delta t$ being time interval (of the time series) between each decomposition step.
@@
The results of this analysis is shown in Fig.5, which shows the full spectrum LE. In the figure, the axes are the three dominant LEs $(\lambda_{1,2,3})$ and the red and blue circles respectively denotes the cases $\rho_{{\rm ext}}>0$ and $<0$. Apart from the fact that all $\lambda_{1,2,3}>0$ which signifies chaos, the accumulation of the red circles at the upper portion of the figure indicate a considerably stronger chaos for $\rho_{{\rm ext}}>0$ than $<0$. The four cases corresponding to four circles in the figure (red and blue) are for different debris velocities $v_{d}=0.2,0.5,1.0,$ and $1.5$ (The numerical code used to calculate the full-spectrum LE is available).
@@
Figure 6: Suppression of density oscillations with periodic charge fluctuation of the negatively charged debris. The top panels (a), (b) are obtained from PIC simulation while the bottom panels (c), (d) are obtained from FCT simulation. The panels on the left are without any charge fluctuation and those on the right are with charge fluctuation.
@@ 2.5 Nonlinear Landau damping?
We now present a combined result of PIC and FCT simulations which points out a phenomenon similar to what is known as “nonlinear Landau damping” [29]. In both cases, we observe a suppression of oscillations when $v_{d}\sim c_{i}$, where $c_{i}=\sqrt{T_{e}/m_{i}}$ is the phase velocity of the ion-acoustic oscillations. However, the fundamental physics behind this observation is slightly different in both cases.
Figure 7: Flattening of the ion VDF due to fluctuation. While the top panels (a) and (b) display respectively the VDFs calculated from the PIC simulation without $(\nu=0)$ and with $(\nu\sim 0.7)$ fluctuation, the bottom panels (c) and (d) show the numerically calculated FWHMs for the corresponding VDFs. The panel (d) on the right clearly shows that the VDF is flattening due to fluctuation.
@@
This effect is shown in Fig.6 for negatively charged debris, both as obtained from PIC and FCT simulations. However, this resonant suppression of oscillation occurs at $v_{d}\lesssim c_{i}$ for PIC simulation and at $v_{d}\gtrsim c_{i}$ for FCT simulation as can be seen from the figure. This discrepancy can be explained from the fact that the PIC simulation captures the full kinetic response including wave-particle resonance and phase space trapping while the FCT simulation captures only the macroscopic fluid-like wave-wave response and lacks phase mixing. In Fig.7, we have shown the ion VDF obtained from PIC simulation data with and without debris charge fluctuation, which clearly show the flattening of the VDF with fluctuation. The suppression of oscillations combined with the definitive proof of flattening of the ion VDF in PIC simulation strongly supports the fact that oscillations undergo NLLD due to charge fluctuation in the case of negatively charged debris. We note here that this flattening of ion VDF is not seen in case of fluctuating positively charged debris.
@@ 3.1 Torus breakdown of a chaotic plasma model
In order to facilitate a dynamical analysis, we make a coordinate transformation to the moving frame of the debris with the introduction of a scaled time $\tau=t-x/v_{d}$, so that in terms of the scaled variable we have $\partial/\partial t\to\partial/\partial\tau$ and $\partial/\partial x\to-v_{d}^{-1}\partial/\partial\tau$ [35]. With this transformation, our plasma model represented by Eqs.(3-6) becomes

### v3

@@ 2.5 Fluid model – FCT simulation
The FCT formalism requires the hydrodynamic equations Eqs.(8-9) to be put in the form of a generalized continuity equation,
$\frac{\partial f}{\partial t}=-\frac{\partial}{\partial x}(fv_{i})+\frac{\partial s}{\partial x},$ (26)
where $f=(n_{i},v_{i})$ is the physical quantity to be solved, $(fv_{i})$ is the corresponding flux, and $s$ is the source term. The external charged debris is entered through the Poisson equation Eq.(6) as before which is to be solved at each time step along with Eq.(26). The external charged debris is modeled with the same profile as used in the PIC simulation, given by Eq.(16).
@@
$\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{m_{\rm FCT}},\quad m_{{\rm FCT}}=6,$ (27)
@@
In the above expression, $R$ is the upper-triangular matrix as obtained from the QR decomposition, $p$ is the index of time step at which the decomposition is carried out, $N$ is total number of such steps, and $T$ is the physical time $T=\sum_{k=1}^{N}\Delta t_{k}$ with $\Delta t$ being the time interval (of the time series) between each decomposition step.
@@
The results of this analysis is shown in Fig.6, which shows the full spectrum LE. In the figure, the axes are the three dominant LEs $(\lambda_{1,2,3})$ and the red and blue circles respectively denote the cases $\rho_{{\rm ext}}>0$ and <0<0. Apart from the fact that all $\lambda_{1,2,3}>0$ which signifies chaos, the accumulation of the red circles at the upper portion of the figure indicate a considerably stronger chaos for $\rho_{{\rm ext}}>0$ than <0<0. The four cases corresponding to four circles in the figure (red and blue) are for different debris velocities $v_{d}=0.2,0.5,1.0$ and $1.5$ (the numerical code used to calculate the full-spectrum LE is available).
@@
Figure 7: Suppression of density oscillations with periodic charge fluctuation of the negatively charged debris. The top panels (a), (b) are obtained from PIC simulation while the bottom panels (c), (d) are obtained from FCT simulation. The panels on the left are without any charge fluctuation and those on the right are with charge fluctuation. The vertical dashed line denotes the position of the debris.
@@ 2.6 Nonlinear Landau damping?
We now present a combined result of PIC and FCT simulations which point out a phenomenon similar to what is known as “nonlinear Landau damping” [29]. In both cases, we observe a suppression of oscillations when $v_{d}\sim c_{i}$, where $c_{i}=\sqrt{T_{e}/m_{i}}$ is the phase velocity of the ion-acoustic oscillations. However, the fundamental physics behind this observation is slightly different in both cases.
Figure 8: Flattening of the ion VDF due to charge fluctuation of the negatively charged debris. While the top panels (a) and (b) display respectively the VDFs calculated from the PIC simulation without $(\nu=0)$ and with $(\nu\sim 0.7)$ fluctuation, the bottom panels (c) and (d) show the numerically calculated FWHMs for the corresponding VDFs. The panel (d) on the right clearly shows that the VDF is flattening due to fluctuation.
@@
This effect is shown in Fig.7 for negatively charged debris, both as obtained from PIC and FCT simulations. However, this resonant suppression of oscillation occurs at $v_{d}\lesssim c_{i}$ for PIC simulation and at $v_{d}\gtrsim c_{i}$ for FCT simulation as can be seen from the figure. This discrepancy can be explained from the fact that the PIC simulation captures the full kinetic response including wave-particle resonance and phase space trapping while the FCT simulation captures only the macroscopic fluid-like wave-wave response and lacks phase mixing. In Fig.8, we have shown the ion VDF obtained from PIC simulation data with and without debris charge fluctuation, which clearly show the flattening of the VDF with fluctuation. The suppression of oscillations combined with the definitive proof of flattening of the ion VDF in PIC simulation strongly supports the fact that oscillations undergo NLLD due to charge fluctuation in the case of negatively charged debris. We note here that this flattening of ion VDF is not seen in case of fluctuating positively charged debris. In the above analysis, we have used $v_{d}=0.8$.
@@ 3.1 Torus breakdown of a chaotic plasma model
In order to facilitate a dynamical analysis, we make a coordinate transformation to the moving frame of the debris with the introduction of a scaled time $\tau=t-x/v_{d}$, so that in terms of the scaled variable we have $\partial/\partial t\to\partial/\partial\tau$ and $\partial/\partial x\to-v_{d}^{-1}\partial/\partial\tau$ [38]. With this transformation, our plasma model represented by Eqs.(8-11) becomes
@@ 3.1.1 Torus breakdown
The reason for the quick transition to a chaotic regime in the case of PIC simulation compared to the analytical model is that the analytical model is a reduced fluid model, which does not have any kinetic effects and provides the dynamics for only a single fluid element. Whereas the PIC model inherently has all the possible kinetic effects such as nonlinear Landau damping, phase-mixing, and particle trapping and nonlinearities making it intrinsically fast and nonlinear. Besides, it shows the cumulative dynamics of millions of particles with many directions for making some of the Lyapunov exponents grow fast.
@@ Acknowledgement
The authors would also like to thank the anonymous referees as their critical comments and suggestion have greatly improved the manuscript.
Appendix A Collisionality and kinetic simulation
It can be argued about the validity of the kinetic simulation in the light of the regime of collisionality. In this appendix, we establish the validity of the kinetic modeling taking consideration of the $e$-$i$ collision by evaluating the Knudsen number. We can derive the Knudsen number (Kn) from the first principle in plasma considering the fact that the dominant collisions are between electrons and ions (this causes the highest momentum transfer). So, by definition, we have
$\textrm{Kn}=\frac{\lambda_{\textrm{mean}}}{L},$ (58)
where $\lambda_{\textrm{mean}}$ is the mean free path for $e$-$i$ collisions and $L$ is the characteristic length of the system. As compared to the ions, electrons are highly mobile (in our case as well), we can assume that
$\lambda_{\textrm{mean}}\simeq\frac{v_{\textrm{th}e}}{\nu_{ei}},$ (59)
where the electron thermal velocity, $v_{\textrm{th}e}=\sqrt{T_{e}/m_{e}}$ and electron-ion collision frequency, $\nu_{ei}$ for Coulomb cross section can be approximated as [35]
$\nu_{ei}\sim\frac{ne^{4}}{16\pi\epsilon_{0}^{2}m_{e}^{2}v_{\textrm{th}e}^{3}},$ (60)
where $n$ is the average plasma density. Our simulation box is of length $L\sim 140\lambda_{D}\sim 1\,\textrm{cm}$. Plugging all the numbers for our simulation parameters, we arrive at the value for the Knudsen number as
$\textrm{Kn}\sim 15\gg 1,$ (61)
which shows that the regime is almost collision-free and the use of kinetic theory is quite justified (through the PIC model).
It can further be argued about the use of the fluid model (Section 2.4) when $\textrm{Kn}\gg 1$. We note that the fluid equations used in this work (or in plasma physics contexts as well) are weighted average of the Boltzmann-Vlasov equation rather than the collisional Boltzmann equation (using Chapman-Enskog expansion). Naturally these fluid equations are valid even in the limit of zero collisions, which is due to the fact that the plasma dynamics is governed by long range Coulomb force rather than Newtonian collisional dynamics (like a hydrodynamic gas).